OpenStudy (kpop4life123):

A set of equations is given below: Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations?

OpenStudy (kpop4life123):

x + 1 = 4x + 5 x = 4x + 5 x + 1 = 4x x + 5 = 4x + 1

OpenStudy (kpop4life123):

the second one x = 4x + 5

OpenStudy (isaidavila):

I'd say the first. Obviusly both equation have something in common y so you have to match both equations

OpenStudy (kpop4life123):

thank you @isaidavila

OpenStudy (briannabug):

But he said he got it right when he clicked on B?